OpenStudy (zzr0ck3r):

how to solve for x? 3^x > x+4

OpenStudy (swissgirl):

well I am a cheater I just graphed the 2 functions

OpenStudy (zzr0ck3r):

hehe, still what is the exact point ?

OpenStudy (swissgirl):

x>1.5619188

OpenStudy (zzr0ck3r):

well first off there will be equal at two points, but I need to know how to do it without calc.

OpenStudy (swissgirl):

ya whoops i needed to find the intersection for the other point. I can give it to u butt i dont know how to solve this without a calc. it gets confusing with ln and stuff

OpenStudy (zzr0ck3r):

yeah I know the two points, but not exact.

OpenStudy (zzr0ck3r):

wolfram does something funny.

OpenStudy (swissgirl):

so \( x<-3.987483 \text { and } x>1.5619188 \)

OpenStudy (zzr0ck3r):

yeh, I dont think this is solvable at our level. http://www.wolframalpha.com/input/?i=solve+3%5Ex+%3E+x%2B4++ check that out

OpenStudy (zzr0ck3r):

yeah I dont know the two base functions they use. maybe next year ill try again:)

OpenStudy (zzr0ck3r):

tnks for looking and thinking:)

OpenStudy (swissgirl):

next yr? what is this for?

OpenStudy (zzr0ck3r):

nothing. I am a math major, so im thinking maybe ill learn about those two functions by then.

OpenStudy (swissgirl):

I am a math major and dont know a thing :(

OpenStudy (zzr0ck3r):

I think the higher we go, the more we dont know.

OpenStudy (zzr0ck3r):

what class you in now?

OpenStudy (swissgirl):

ummm its more of a advanced discrete math class. Just alot of proving

OpenStudy (swissgirl):

ok dumbcow will have the answer lol

OpenStudy (zzr0ck3r):

yah im taking descrete in the fall. Was actually helping a girl today in descrete prove that something was one-to-one. (we covered it in some proof writing class)

OpenStudy (dumbcow):

yeah, these type of problems can't be solved algebraically ..no i can't do it either, but thats why there is solving analytically with technology :)

OpenStudy (dumbcow):

btw i was math major as well and i never was taught the product log function that wolfram uses

OpenStudy (swissgirl):

hahah i still feel stupid cant prove a dumb thing

OpenStudy (zzr0ck3r):

ahh ty for info.

OpenStudy (zzr0ck3r):

what dumb thing?

OpenStudy (swissgirl):

i cant prove nething at all. Like I always get stuck at some point or other

OpenStudy (zzr0ck3r):

well it would not be fun any other way:)

OpenStudy (swissgirl):

hahahah true I just learn it and never have time to digest nething. I am always running behind. Wish there was more time on my hands

OpenStudy (dumbcow):

if you had to this by hand without graphing software...i would approximate using Newtons method for finding zeros \[x_{n+1} =x_{n}-\frac{f(x_{n})}{f'(x_{n})} \] Let f(x) = 3^x -x-4 , then f'(x) = 3^x ln(3) - 1